OpenStudy (anonymous):

The area of a triangle is 10 square inches and its height is 5 inches. Find the base. 4 in. 25 in.

OpenStudy (kamibug):

The formula for the area of a triangle is: A= bh / 2 To solve for the base, we can just substitute in what we already know. 10 = b5 / 2 Solve this equation like you would any other. 10 = b5 /2 20 = b5 4 = b Then check your work... :) 10 = 5(4) / 2 10 = 20 / 2 10 = 10 So the base is 4 inches. :) Hope this helped!

OpenStudy (anonymous):

thank u